alextrejo373:

Circle A has a diameter of 8 inches, a circumference of 25.12 inches, and an area of 50.24 square inches. The diameter of circle B is 3 inches, the circumference is 9.42 inches, and the area is 7.065 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)​

HeartTricker:

Hi, it says "Using the formula for circumference, " Do you know the formula for find the circumference of a circle ?

alextrejo373:

multiply the diameter of the circle with π (pi). The circumference can also be calculated by multiplying 2×radius with pi (π=3.14).

HeartTricker:

correct but its can aslo be \[ C= 2 \pi r \] we use the Greek letter π because if we tried to write out the actual number, it goes on forever.

HeartTricker:

now they say Circle A has a diameter of 8 inches what is the radius of circle A? any idea ?

alextrejo373:

r=4in d Diameter =8 in

HeartTricker:

pi * 8 = 25.12 now "solve for pi" by dividing both sides by 8 pi= 24.12/8 what do you get?

alextrejo373:

Part A Circle A Circumference = 2πr 25.12 = 2π(4) 25.12 = 8π Divide both side by 8 3.14 = π Circle B Circumference = 2πr 9.42 = 2π(3/2) 9.42 = 3π Divide both side by 3 3.14 = π Part B Circle A A = πr² 50.24 = π(4)² 50.24 = 16π Divide both side by 16 3.14 = π Circle B A = πr² 7.065 = π(3/2)² 7.065 = 9π/4 Divide both side by 9/4 3.14 = π Part C He/She used exactly 3.14 for the value of pi in circle A and B to get the circumference and the area.